OpenStudy (anonymous):

(5c+2/3)(5c-2/3) foil method. i know step one would be 5c* 5c = c^25 can you help me with next steps?

OpenStudy (anonymous):

do you know the distribution property?

OpenStudy (anonymous):

My teacher said to use the foil method? Is that what you are talking about?

OpenStudy (anonymous):

or should i say the "distributive" property of multiplication over addition. a(b+c)=ab+ac

OpenStudy (anonymous):

no I'm sorry can you explain?

OpenStudy (anonymous):

3(2+4)= 3*6=18=6+12=3*2+3*4 the distributive property let's you skip the middle stuff, especially when you can't perform the addition in tha parentheses. ex. 3(x+5) = 3x +15 have you seen things like this before?

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so in my problem I would first do what? take the 5c and the 5c right which is c^25?

OpenStudy (anonymous):

that's all that foil is... except you do it twice. you have \[\left(5c+\frac{2}{3}\right)\left(5c-\frac{2}{3}\right)\]pretend \[\left(5c+\frac{2}{3}\right)=a\]then you have \[a\left(5c-\frac{2}{3}\right)=a(5c)-a\frac{2}{3}\]now just put \[\left(5c+\frac{2}{3}\right)\]back in for a and distribute again\[\left(5c+\frac{2}{3}\right)\cdot5c-\left(5c+\frac{2}{3}\right)\cdot\frac{2}{3}=(5c)^2+(5c)\cdot\frac{2}{3}-(5c)\cdot\frac{2}{3}-\left(\frac{2}{3}\right)^2\]work that out and you're done

OpenStudy (anonymous):

I keep coming up with a fraction is that right?

OpenStudy (anonymous):

25 c^2-4\/9

OpenStudy (anonymous):

that's the right answer!